# Notes on the Practical Application of Nested Sampling: MultiNest, (non)convergence, and rectification

### v1

@@ Abstract
Nested sampling is a promising tool for Bayesian statistical analysis because it simultaneously performs parameter estimation and facilitates model comparison. MultiNest is one of the most popular nested sampling implementations, and has been applied to a wide variety of problems in the physical sciences. However, MultiNest results are frequently unreliable, and accompanying convergence tests are a necessary component of any analysis. Using simple, analytically tractable test problems, I illustrate how MultiNest (1) can produce systematically biased estimates of the Bayesian evidence, which are more significantly biased for problems of higher dimensionality; (2) can derive posterior estimates with errors on the order of $\sim 100\%$; (3) is more likely to underestimate the width of a credible interval than to overestimate it — to a minor degree for smooth problems, but much more so when sampling noisy likelihoods. Nevertheless, I show how MultiNest can be used to jump-start Markov chain Monte Carlo sampling or more rigorous nested sampling techniques, potentially accelerating more trustworthy measurements of posterior distributions and Bayesian evidences, and overcoming the challenge of Markov chain Monte Carlo initialization.
@@ 1 Introduction
MultiNest (Feroz & Hobson, 2008; Feroz et al., 2009) is arguably the most popular nested sampling implementation. It is relatively fast (see, for example, Figure 7 of Handley et al. (2015) or Figure 4 of Albert (2020)), and it can provide accurate results for simple problems with few parameters. MultiNest has been applied to X-ray astronomy (e.g., Buchner et al., 2014; Miller et al., 2019; Riley et al., 2019), the study of exoplanets (e.g. Kipping, 2013; Morton et al., 2016; Tsiaras et al., 2018), and cosmology (e.g. Planck Collaboration et al., 2014, 2016, 2020), in addition to many other fields.
@@
However, numerous studies have found that MultiNest can perform poorly and unreliably, both in terms of inferring posterior distributions (e.g., Miller et al., 2019; Ih & Kempton, 2021; Miller et al., 2021) and measuring Bayesian evidences (e.g., Albert, 2020; Nelson et al., 2020; Lemos et al., 2023). One of the challenges faced when using MultiNest is deciding on values for its free parameters. The most important of these parameters are the number of live points ($N_{\rm live},$ which determines how densely parameter space is sampled) and the sampling efficiency ($\epsilon$, which controls how conservatively the algorithm explores parameter space). One can also tune the algorithm’s stopping criterion through a tolerance parameter for the evidence (tol), but that is usually of little consequence to the algorithm’s performance. [^1] Unless one’s goal is maximum-likelihood estimation: As MultiNest marches higher and higher toward the peak of the likelihood function, eventually the remaining prior volume is so minuscule that the remainder of the evidence integral is negligible, and the algorithm terminates. Setting a smaller tolerance value continues iteration until less and less of the evidence integral remains.
@@
Unfortunately, the values for these parameters suggested by the MultiNest developers are often inadequate (see Section 4). For example, the README file included with the MultiNest source code suggests using $\epsilon=0.8$ for parameter inference and $\epsilon=0.3$ for evidence measurement. [^2] https://github.com/farhanferoz/MultiNest/blob/cc616c17ed440fbff8d35937997c86256ed4d744/README.md The primary MultiNest paper states that “we found 400 and 50 active points to be adequate for evidence evaluation and parameter estimation respectively,” and suggests sampling efficiency values of $1$ and $0.3$ for parameter estimation and evidence evaluation (Feroz et al., 2009, Section 8). As shown in Section 4, these suggestions about sampling posteriors are actually reasonable for simple Gaussian distributions (as long as one can accept a $\sim 10\%$ error for each posterior distribution). However, even for ten dimensional Gaussian distributions these suggested parameter values are unable to accurately measure the Bayesian evidence. [^3] It is worth noting that the authors of MultiNest have advocated that “it is always advisable to increase $N_{\rm live},$ and reduce [$\epsilon$] to check if the posteriors and evidence values are stable,” (Feroz et al., 2019) although many papers have used MultiNest without reporting such convergence tests.
@@
Many studies have investigated potential deficiencies in nested sampling algorithms, developed algorithmic improvements, and suggested methods for testing convergence. To give a few examples, Nelson et al. (2020); Albert (2020); Lemos et al. (2023) have noted systematic, and sometimes dimension- and sampling efficiency-dependent, biases in the evidence estimates produced by MultiNest. Buchner (2016a) developed a test that illustrated MultiNest’s tendency to sample from an erroneously small prior volume, and developed an alternative algorithm for unbiased, albeit slower, sampling. Higson et al. (2019a) proposed some strategies to check convergence by cross-checking multiple analyses, and advocated that when using MultiNest one should decrease $\epsilon$ and increase $N_{\rm live}$ to check convergence.
@@
My aims here are to systematically study MultiNest’s convergence in terms of both the sampling efficiency and number of live points for problems of varying dimensionality, and to present strategies that can be applied to improve upon completed nested sampling analyses. Section 2 provides a brief overview of MultiNest and nested sampling, after which Section 3 describes some of my methodology. The results of my tests using MultiNest to sample smooth distributions are presented in Section 4, and my exploration of MultiNest’s performance on a noisy target distribution is presented in Section 5. Section 6 demonstrates a couple of methods that can be used to derive better posterior and evidence estimates, using an initial MultiNest (or other nested sampling) run to accelerate the process.
@@ 2 The MultiNest Algorithm
The sampling efficiency factor $\epsilon$ acts as a safeguard in the algorithm, making it more likely that new points are actually drawn from within the intended likelihood contour. Unfortunately, when using $\epsilon=0.3$, for example, MultiNest is still wont to sample from too small a volume of parameter space (Buchner, 2016a). Notably, the amount of ellipsoid expansion in any given linear dimension at constant $\epsilon$ is much smaller in higher dimensions. This is why smaller efficiencies are required to produce unbiased results in higher dimensions and why using sampling efficiencies greater than unity is much more damaging for lower-dimension problems, as shown in Section 4.
@@ 3 Test Problems and Methodology
The one-dimensional posterior distributions in the normal and log-Gamma distribution tests are trivial to calculate, so one can easily test the ability of each sampler to correctly recover the underlying posterior distribution. I took a practical approach to estimating errors in posterior recovery: I first computed a kernel-density estimate using the posterior samples drawn in each test, and then calculated the $L_{1}$ error of that kernel density estimate and the true underlying distribution. Explicitly, I calculated in each dimension ($i$)
@@
such that positive values of $\Delta\log{Z}$ indicate that MultiNest overestimated the evidence.
@@ 3.1.1 The Normal Distribution
One would hope that any Bayesian analysis tool is capable of sampling this distribution adequately. In practice, however, unless used with sufficiently rigorous “resolution settings,” MultiNest fails to recover the correct evidence in 10- or 20-dimensional test problems. For each problem, I selected standard deviations $\sigma_{i}$ uniformly between 0.1 and 0.5 (inclusive), with a spacing dependent on the number of dimensions.
@@ 3.1.2 An Extended Rosenbrock Distribution
I chose $\mu=0$, $a=2.0$, and each $b_{j,i}$ was chosen between 1 and 10 (inclusive) with a geometrically uniform spacing depending on the total problem dimensionality.
@@ 3.1.3 The Log-Gamma Distribution
where $\Gamma$ is the Gamma function, $c=1$, $\mu=2/3$, and scale parameters $\sigma_{i}$ were spaced log-uniformly between 0.01 and 0.05 (inclusive).
@@ 3.1.4 A Noisy Normal Distribution
While attempting to understand why MultiNest might perform fairly reliable posterior inference for many standard test problems but systematically underestimate credible intervals to noticeable degrees when applied to astronomical data analyses (e.g., Miller et al., 2019; Ih & Kempton, 2021; Miller et al., 2021), I considered the possibility that some sort of “noise” (perhaps some fairly high-frequency or stochastic variation in the likelihood as a function of one or more parameters, superimposed over some smooth trend) might be present in actual analyses but not the smooth test problems described above. To test this idea, I used the following log-likelihood function in N dimensions
@@
where $f(x)$ is a linear interpolant over a sequence of pseudorandom numbers drawn uniformly between -1 and 1, and $A$ is a noise amplitude. For simplicity I used the same pseudorandom field in each dimension, which had between 250 and 1250 samples per $\sigma_{i}$. To be clear, I make no claim that this sort of noise is actually present in practical analyses: this is simply one test case where MultiNest performs similarly to the aforementioned applications to real data.
@@ 3.2.1 Smooth Test Problems
I have focused on only the standard MultiNest algorithm (Feroz et al., 2009, version 3.12), without using importance nested sampling (Feroz et al., 2019). Since this variation does not alter MultiNest’s parameter exploration strategy, it would not ameliorate any of the cases where MultiNest’s posterior estimates severely erred. Importance nested sampling might improve performance in some of the cases where the posterior estimates were acceptable and the evidence was simply overestimated, although it is far from panacean (e.g., Nelson et al., 2020), sometimes leading to the confident exclusion of the true evidence (Buchner, 2016a).
Although each problem was unimodal, I allowed MultiNest to perform mode separation if it identified multiple modes, since in practice one rarely knows a priori how many modes are present in a given problem. I performed a handful of log-gamma tests with mode separation deactivated and observed no significant changes in the results. I set MultiNest’s evidence tolerance parameter to $10^{-5}$ in each analysis to avoid the possibility that parameter space might not be thoroughly explored because the tolerance was too large, although I also repeated $\sim 100$ normal and log-gamma tests using a tolerance parameter of $0.25$ and found no appreciable differences in the quality of the results.
@@ 3.2.2 A Noisy Test Problem
For the test introduced in Section 3.1.4, developed in an attempt to understand a possible cause for MultiNest systematically underestimating credible region widths (e.g., Miller et al., 2019; Ih & Kempton, 2021) in a sampling-efficiency-dependent way (e.g., Miller et al., 2021), [^7] Most quantitatively, in Miller et al. (2021), MultiNest underestimated the neutron star equatorial radius posterior width by a factor of $\sim 0.67$ when used with ($N_{\rm live}=1000,$ $\epsilon=0.01$), a factor of $\sim 0.79$ when used with ($N_{\rm live}=3000,$ $\epsilon=0.01$), and even smaller factors when used with larger sampling efficiencies. I focused instead on MultiNest parameters closer to those used in scientific applications. Specifically, I used $N_{\rm live}\in\{300,1000,3000,10000\}$, for each $N_{\rm live}$ I have used the sampling efficiencies $\epsilon~{}\in~{}\{0.3,0.1,0.03,0.01\}$, and I have tested each ($N_{\rm live}$, $\epsilon$) combination 10 independent times. For each of these, I have saved MultiNest outputs at two evidence tolerance values, $\texttt{tol}\in\{0.5,0.01\}$, waiting for the first series to complete and then re-starting each with a lower tolerance. For succinctness, I have only reported in Section 5 results from a 10D test case.
@@ 4.1 Normal Distributions
  • MultiNest is slightly biased towards underestimating credible region widths.
@@
Figure 3 displays the errors in the width of the ±$\pm 1\sigma$ credible intervals for each of the normal distribution analyses, collated similarly to the $L_{1}$ error measurements. Out of the 15750 posterior distributions inferred in these tests, MultiNest produced an erroneously narrow posterior distribution on 8370 instances. Although MultiNest only underestimated the width of posterior credible regions $\sim 53\%$ of the time, the odds of this many or more low estimates occurring in a sample of unbiased posterior estimates (that is, with widths overestimated as often as they are underestimated) are on the order of $\sim 10^{-13}\%$. Restricting the results to the 15477 posterior inferences that had $L_{1}$ errors less than 0.2, excluding the cases where MultiNest estimated the posteriors particularly poorly, MultiNest underestimated the width of the credible region in 8168 cases. With an unbiased sampler the odds of getting that many (or more) low rolls is a much more reasonable $\sim 2\times 10^{-12}\%$. Although not by particularly harmful amounts for this simple problem, MultiNest appears to be systematically biased towards reporting erroneously narrow credible regions.

### v2

@@ Abstract
Nested sampling is a promising tool for Bayesian statistical analysis because it simultaneously performs parameter estimation and facilitates model comparison. MultiNest is one of the most popular nested sampling implementations, and has been applied to a wide variety of problems in the physical sciences. However, MultiNest results, like those of any sampling tool, can be unreliable, and accompanying convergence tests are a necessary component of any analysis. Using analytically tractable test problems, I illustrate how MultiNest, when applied without rigorously chosen hyperparameters, (1) can produce systematically erroneous estimates of the Bayesian evidence, which are more significantly biased for problems of higher dimensionality; (2) can derive posterior estimates with errors on the order of $\sim 100\%$; (3) can, particularly when sampling noisy likelihood functions, systematically underestimate posterior widths. Furthermore, I show how MultiNest, thanks to the advantageous speed at which it explores parameter space, can also be used to jump-start Markov chain Monte Carlo sampling or more rigorous nested sampling techniques, potentially accelerating more robust measurements of posterior distributions and Bayesian evidences, and overcoming the challenge of Markov chain Monte Carlo initialization.
@@ 1 Introduction
MultiNest (Feroz & Hobson, 2008; Feroz et al., 2009) is arguably the most popular nested sampling implementation. It is relatively fast (see, for example, Figure 7 of Handley et al. (2015) or Figure 4 of Albert (2020)), and it can provide accurate results when used with sufficiently rigorous settings. MultiNest has been applied to X-ray astronomy (e.g., Buchner et al., 2014; Miller et al., 2019; Riley et al., 2019), the study of exoplanets (e.g. Kipping, 2013; Morton et al., 2016; Tsiaras et al., 2018), and cosmology (e.g. Planck Collaboration et al., 2014, 2016, 2020), in addition to many other fields.
@@
However, numerous studies have found that MultiNest can perform poorly and unreliably, both in terms of inferring posterior distributions (e.g., Miller et al., 2019; Ih & Kempton, 2021; Miller et al., 2021) and measuring Bayesian evidences (e.g., Albert, 2020; Nelson et al., 2020; Lemos et al., 2023). One of the challenges faced when using MultiNest, like virtually all Bayesian analysis tools, is deciding on values for its free parameters. The most important of these parameters are the number of live points ($N_{\rm live},$ which determines how densely parameter space is sampled) and the sampling efficiency ($\epsilon$, which controls how conservatively the algorithm explores parameter space). One can also tune the algorithm’s stopping criterion through a tolerance parameter for the evidence (tol), but that is usually of little consequence to the algorithm’s performance. [^1] Unless one’s goal is maximum-likelihood estimation: As MultiNest marches higher and higher toward the peak of the likelihood function, eventually the remaining prior volume is so minuscule that the remainder of the evidence integral is negligible, and the algorithm terminates. Setting a smaller tolerance value continues iteration until less and less of the evidence integral remains. Regardless, the small number of free parameters in the MultiNest algorithm are an attractive feature, making testing its convergence a straightforward process.
@@
Unfortunately, the values for these parameters suggested by the MultiNest papers and documentation, upon which many users rely, are often inadequate (see Section 4). For example, the README file included with the MultiNest source code suggests using $\epsilon=0.8$ for parameter inference and $\epsilon=0.3$ for evidence measurement. [^2] https://github.com/farhanferoz/MultiNest/blob/cc616c17ed440fbff8d35937997c86256ed4d744/README.md The primary MultiNest paper states that “we found 400 and 50 active points to be adequate for evidence evaluation and parameter estimation respectively,” and suggests sampling efficiency values of $1$ and $0.3$ for parameter estimation and evidence evaluation (Feroz et al., 2009, Section 8). As shown in Section 4, these suggestions about sampling posteriors are reasonable for simple Gaussian distributions. However, even for ten dimensional Gaussian distributions these suggested parameter values are inadequate to accurately measure the Bayesian evidence. [^3] It is worth noting that the authors of MultiNest have advocated that “it is always advisable to increase $N_{\rm live},$ and reduce [$\epsilon$] to check if the posteriors and evidence values are stable,” (Feroz et al., 2019) although many papers have used MultiNest without reporting such convergence tests.
@@
Many studies have investigated potential deficiencies in nested sampling algorithms, developed algorithmic improvements, and suggested methods for testing convergence. To give a few examples, Nelson et al. (2020); Albert (2020); Lemos et al. (2023) have noted systematic, and sometimes dimension- and sampling efficiency-dependent biases in the evidence estimates produced by MultiNest. Buchner (2016a) developed a test that illustrated MultiNest’s tendency to sample from an erroneously small prior volume, and developed an alternative algorithm for unbiased, albeit slower, sampling. Higson et al. (2019a) proposed some strategies to check convergence by cross-checking multiple analyses, and advocated that when using MultiNest one should decrease $\epsilon$ and increase $N_{\rm live}$ to check convergence.
@@
My aims here are to systematically study MultiNest’s convergence in terms of both the sampling efficiency and number of live points for problems of varying dimensionality, and to present strategies that can be applied to improve upon completed nested sampling analyses , taking advantage of MultiNest’s favorable speed. Section 2 provides a brief overview of MultiNest and nested sampling, after which Section 3 describes some of my methodology. The results of my tests using MultiNest to sample smooth distributions are presented in Section 4, and my exploration of MultiNest’s performance on a noisy target distribution is presented in Section 5. Section 6 demonstrates a couple of methods that can be used to derive better posterior and evidence estimates, using an initial MultiNest (or other nested sampling) run to accelerate the process.
@@ 2 The MultiNest Algorithm
The sampling efficiency factor $\epsilon$ acts as a safeguard in the algorithm, making it more likely that new points are actually drawn from within the intended likelihood contour. However, when using $\epsilon=0.3$, for example, MultiNest is still wont to sample from too small a volume of parameter space (Buchner, 2016a). Notably, the amount of ellipsoid expansion in any given linear dimension at constant $\epsilon$ is much smaller in higher dimensions. This is why smaller efficiencies are required to produce accurate results in higher dimensions and why using sampling efficiencies greater than unity is much more damaging for lower-dimension problems, as shown in Section 4.
@@ 3 Test Problems and Methodology
The one-dimensional posterior distributions in the normal and log-Gamma distribution tests are trivial to calculate, so one can easily test the ability of each sampler to correctly recover the underlying posterior distribution. I took a practical approach to estimating errors in posterior recovery: I first computed a kernel-density estimate using the posterior samples drawn in each test (between $\sim 2000$ and $\sim 530000$ samples depending predominantly on the problem dimensionality and number of live points), and then calculated the $L_{1}$ error of that kernel density estimate and the true underlying distribution. Explicitly, I calculated in each dimension ($i$)
@@
such that positive values of $\Delta\log{Z}$ indicate that MultiNest overestimated the evidence. This sort of evidence error analysis has been conducted previously in Brewer et al. (2011); Albert (2020); Nelson et al. (2020); Lemos et al. (2023), for example.
@@ 3.1.1 The Normal Distribution
For each problem, I selected standard deviations $\sigma_{i}$ uniformly between 0.1 and 0.5 (inclusive), with a spacing dependent on the number of dimensions. For this problem, I used uniform priors in each dimension on the interval $[-20,20]$. Because this problem is unimodal with a likelihood gradient only increasing further from its peak, the entire prior serving as an effective basin of attraction, the prior widths have negligible influence, and do not affect any of the conclusions drawn from these tests. [^7] A quick test one can run to verify this is sampling a 20-dimensional normal distribution, as described here, using $N_{\rm live}=300$ and $\epsilon=0.3$. As an example, using the default priors adopted here, ten independent analyses had a mean log evidence error of $\Delta\log{Z}=1.35$ with a standard deviation of $0.26$. Shrinking the priors to $[-5,5]$ in each dimension and running again, I found a mean error of $\Delta\log{Z}=1.20$ with a standard deviation of $0.31$. Similar results hold with more live points (see section 4.1) or even tighter priors.
@@ 3.1.2 An Extended Rosenbrock Distribution
I chose $\mu=0$, $a=2.0$, and each $b_{j,i}$ was chosen between 1 and 10 (inclusive) with a geometrically uniform spacing depending on the total problem dimensionality. For this problem, I used uniform priors in each dimension on the interval $[-50,50]$. As with the normal distribution, the priors are largely irrelevant, provided they are wide enough for the analytical evidence value to apply.
@@ 3.1.3 The Log-Gamma Distribution
where $\Gamma$ is the Gamma function, $c=1$, $\mu=2/3$, and scale parameters $\sigma_{i}$ were spaced log-uniformly between 0.01 and 0.05 (inclusive). For this problem, I used uniform priors in each dimension on the interval $[-10,2]$, and unlike the other tests, in this case priors can have a substantial effect. As discussed in Section 4.3, MultiNest has a tendency to miss the peak of this distribution and instead hone in on some portion of its long tail in various numbers of dimensions. More narrow priors, particularly on the lower end, would limit the range over which MultiNest could get stuck; wider or narrower priors can thus lead to greater or lower magnitude evidence errors, although the priors have little-to-no effect on which hyperparameters are necessary to thoroughly sample the distribution.
@@ 3.1.4 A Noisy Normal Distribution
While attempting to understand why MultiNest might perform fairly reliable posterior inference for many standard test problems but in certain astronomical data analyses systematically underestimate credible region widths (e.g., Miller et al., 2019; Ih & Kempton, 2021; Miller et al., 2021), [^8] Most quantitatively, in Miller et al. (2021), MultiNest underestimated the neutron star equatorial radius posterior width by a factor of $\sim 0.67$ when used with ($N_{\rm live}=1000,$ $\epsilon=0.01$), a factor of $\sim 0.79$ when used with ($N_{\rm live}=3000,$ $\epsilon=0.01$), and even smaller factors when used with larger sampling efficiencies. I considered the admittedly speculative possibility that some sort of “noise” (perhaps some fairly high-frequency or stochastic variation in the likelihood as a function of one or more parameters, superimposed over some smooth trend) might be present in actual analyses but not the smooth test problems described above.999 I was led to this line of investigation by first using a KDE of the posterior samples from Miller et al. (2021) as a target likelihood function and finding that MultiNest produced accurate posterior inferences (for a range of KDE bandwidths), suggesting that some property of the analysis besides simply the shape of the posteriors was to blame. To test this idea, I used the following log-likelihood function in N dimensions
@@
where $f(x)$ is a linear interpolant over a sequence of pseudorandom numbers drawn uniformly between -1 and 1, and $A$ is a noise amplitude. For simplicity I used the same pseudorandom field in each dimension, which had between 250 and 1250 samples per $\sigma_{i}$. The standard deviations and priors were the same as in Section 3.1.1. To be clear, I make no claim that this sort of noise is actually present in practical analyses: this is simply one test case where MultiNest performs similarly to the aforementioned applications to real data.
@@ 3.2.1 Smooth Test Problems
I have focused on only the standard MultiNest algorithm (Feroz et al., 2009, version 3.12), without using importance nested sampling (Feroz et al., 2019). Since this variation does not alter MultiNest’s parameter exploration strategy, it would not ameliorate any of the cases where MultiNest’s posterior estimates erred. Importance nested sampling might improve performance in some of the cases where the posterior estimates were acceptable and the evidence was simply overestimated, although it is far from panacean (e.g., Nelson et al., 2020), sometimes leading to the confident exclusion of the true evidence (Buchner, 2016a).
Although each problem was unimodal, I allowed MultiNest to perform mode separation if it identified multiple modes, since in practice one rarely knows a priori how many modes are present in a given problem. I performed a handful of log-gamma tests with mode separation deactivated and observed no significant changes in the results. I set MultiNest’s evidence tolerance parameter to $10^{-5}$ in each analysis to avoid the possibility that parameter space might not be thoroughly explored because the tolerance was too large, although I also repeated $\sim 100$ normal and log-gamma tests using a tolerance parameter of $0.25$ and found no appreciable differences in the quality of the results. Because each problem is unimodal and none are characterized by log-likelihood plateaus, MultiNest’s performance is largely insensitive to the prior bounds, as described above. However, if multiple modes may be present, as is generally the case, one must be careful to use enough live points for MultiNest to have a good chance of finding each relevant mode within the prior volume (see, for example the discussion in Section 9 of Feroz & Hobson, 2008).
@@ 3.2.2 A Noisy Test Problem
For the test introduced in Section 3.1.4, developed in an attempt to understand a possible cause for MultiNest systematically underestimating credible region widths (e.g., Miller et al., 2019; Ih & Kempton, 2021) in a sampling-efficiency-dependent way (e.g., Miller et al., 2021), I focused instead on MultiNest parameters closer to those used in scientific applications. Specifically, I used $N_{\rm live}\in\{300,1000,3000,10000\}$, for each $N_{\rm live}$ I have used the sampling efficiencies $\epsilon~{}\in~{}\{0.3,0.1,0.03,0.01\}$, and tested each ($N_{\rm live}$, $\epsilon$) combination 10 independent times. For each of these, I have saved MultiNest outputs at two evidence tolerance values, $\texttt{tol}\in\{0.5,0.01\}$, waiting for the first series to complete and then re-starting each with a lower tolerance. For succinctness, I have only reported in Section 5 results from a 10D test case.
@@ 4.1 Normal Distributions
Figure 3 displays the errors in the width of the ±$\pm 1\sigma$ credible intervals for each of the normal distribution analyses, collated similarly to the $L_{1}$ error measurements. For this simple problem, errors in the posterior widths derived by MultiNest behave similarly to their $L_{1}$ errors, shrinking roughly according to $N_{\rm live}^{-1/2}$, and are largely independent of the sampling efficiency provided that $\epsilon\leq 1$; however as shown in Section 5, this independence is not general.
In this test problem isolikelihood surfaces can be perfectly described by hyperellipsoids, which should minimize any errors that could potentially follow from the specific choice of using ellipsoids to construct bounding regions for drawing new points, which should closely match the isolikelihood surface corresponding to the lowest-likelihood live point in order to prevent errors. However, depending on the precise distributions of live points, each bounding ellipsoid can still fail to encompass the intended isolikelihood surface (see, for example, Figure 1 of Lemos et al., 2023), so that samples may be drawn from an artificially small region of the prior volume at each stage. Naturally, this potential problem becomes less significant when more live points and lower sampling efficiencies are used.101010 A similar bias can affect slice samplers if too few random walk iterations are used and freshly drawn points are correlated with previous points. Reassuringly, these errors seem to predominantly affect evidence measurements, and in most cases the posterior errors were independent of the sampling efficiency. The only cases for this problem where MultiNest produced anomalously erroneous posteriors were 5D tests using sampling efficiencies greater than unity. In these cases, MultiNest was forced to draw samples from within an artificially small hypervolume; although the same is true for the higher-dimensionality test with the same efficiency settings, in higher dimensions the shrinkage and expansion in any given dimension is diluted.